OpenStudy (anonymous):

Hi I need to find the solution set for the following equation √(10x+5) - √(6x-11) =6

OpenStudy (anonymous):

you sure that is a "-" and not a "+"?

OpenStudy (anonymous):

Yes the equation is correct

OpenStudy (anonymous):

too bad because if it was a +, 2 would work by inspection. this problem is a pain. you have to square twice

OpenStudy (anonymous):

And it says solve the equation. The solution set is blank and blank

OpenStudy (anonymous):

ok let me work it out. i was trying to guess what would give a perfect square for 10x+5 and 6x-11

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

2 works because 2*10+5=25 and 6*2-11=1

OpenStudy (anonymous):

and 5+1=6

OpenStudy (anonymous):

So would the sol set be 5,1

OpenStudy (anonymous):

no the solution would be x = 2, but it is not. i am still working.

OpenStudy (anonymous):

have to square both sides, then square again. i will show you but i want to get a correct answer so i don't blow it.

OpenStudy (anonymous):

good heavens. i will be back in ten minutes with the answer and the algebra. hold on.

OpenStudy (anonymous):

K :-)

OpenStudy (anonymous):

62

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

would you like a worked out solution?

OpenStudy (anonymous):

square both sided. you will get \[16x-6-2\sqrt{(10x+5)(6x-11)}=36\]

OpenStudy (anonymous):

\[-2\sqrt{(10x+5)(6x-11)}=42-16x\]

OpenStudy (anonymous):

divide by -2 \[\sqrt{(10x+5)(6x-11)}=8x-21\]

OpenStudy (anonymous):

so is there only one solution of 62

OpenStudy (anonymous):

sorry my computer is slow

OpenStudy (anonymous):

then square again. you will get \[(10x-5)(6x-11)=(8x+21)^2\]

OpenStudy (anonymous):

yes only one answer. you will solve a quadratic and get two solutions but only one will work.

OpenStudy (anonymous):

introduced a false solution when you squared

OpenStudy (anonymous):

typo on last answer. it should have been \[(10x+5)(6x-11)=(8x-21)^2\]

OpenStudy (anonymous):

thx u

OpenStudy (anonymous):

expand these to get \[60x^2-80x-55=64x^2-336x+441\]

OpenStudy (anonymous):

set =0 and solve: \[4x^2-256x+496=0\] divide by 4 \[x^2-64x+124=0\]

OpenStudy (anonymous):

factor \[(x-62)(x-2)=0\] \[x=62\] or \[x=2\]

OpenStudy (anonymous):

but there is only one answer because 2 does not work. if you replace x by 2 in the original equation you get \[\sqrt{25}-\sqrt{1}=6\] \[5-1=6\] which is false.

OpenStudy (anonymous):

negatives on both?? thx

OpenStudy (anonymous):

62 will work but only 62, not 2.

OpenStudy (anonymous):

oh negatives on both when factoring. but the only answer is x = 62

OpenStudy (anonymous):

you have to plug it into the original equation

OpenStudy (anonymous):

\[\sqrt{10\times 62+5}-\sqrt{6\times 62-11}\] \[\sqrt{625}-\sqrt{361}=25-19=6\]

OpenStudy (anonymous):

i am confused i just need the solution set

OpenStudy (anonymous):

solution is 62

OpenStudy (anonymous):

62 works, 2 does not

OpenStudy (anonymous):

ok...thx